# Galactic foreground emissions randomization due to chaotic/turbulent dynamics of magnetized plasma dominated by magnetic helicity

### v1

@@ 4 Observations in the Galactic magnetized plasma
In the paper (Hutschenreuter et al. 8) a disentangling of an all-sky map for average intragalactic $B_{LoS}$ from the Faraday effect map was reported. Some complementary tracers of the $n_{e}$ (the free-free map of the Planck survey 31, extra-Galactic Faraday data 42, a $H-\alpha$ map 3, pulsar data 21) were also used for the purpose. and an all-sky map of the electron dispersion measure (the integrated electron density) was also constructed.
Figure 7 shows the power spectrum of $B_{LoS}$. The spectral data were taken from Figs. 9a of the paper (Hutschenreuter et al. 8). The dashed curve in Fig. 7 indicates the best fit by the stretched exponential spectral law Eq. (22).
An all-sky map of the integrated electron density (electron dispersion measure) was also obtained by this method. Figure 8 shows the power spectrum of the electron dispersion measure. The spectral data were taken from Figs. 9b of the paper (Hutschenreuter et al. 8). The dashed curve in Figs. 8 indicates the best fit by the same stretched exponential spectral law Eq. (22). Apparently the magnetic field imposes its degree of randomization (the $\beta=1/4$) on the electron dispersion measure.
@@ 5 Synchrotron emission
The astroparticles (relativistic electrons and positrons) produce the synchrotron emission moving through the magnetic field. Therefore the diffuse (polarized) synchrotron emission is an effective tracer of the magnetic field in the magnetized non-thermal plasma. This emission is also one of the main components of the polarized foreground for the cosmic microwave background radiation. Therefore, studying its spectral properties is important for understanding the magnetized intragalactic plasma’s physical processes and obtaining clean CMB maps.
@@
A spin-2 decomposition of the polarization tensor was suggested in a paper [44]. The polarization tensor is decomposed into two rotationally invariant quantities: scalar E and pseudo-scalar B. The B-mode of this decomposition has magnetic field-type parity.
In a paper [5] results obtained using this technique for the Parkes 2.4 GHz survey of synchrotron polarised emission for the Southern Galactic plane were reported. Figure 10 shows the power spectrum for the B-mode. The spectral data were taken from Fig. 3 of the Ref. (Giardino et al. 5). The dashed curve in Fig. 10 indicates the best fit by the stretched exponential spectral law Eq. (22).
@@ 6 Dust emission
Polarized dust emission is not only one of the most reach sources of information about the Galactic magnetic field (see, for instance, a recent paper by 7, and references therein) but also the main obstacle for observing (detecting) the gravitational waves in primordial CMB B-modes, especially for the high-frequency Planck’s channels [14].
@@
Figure 12: As in Fig 11. but for the T-mode of the thermal polarized dust emission
Figure 13: As in Fig 11. but for the B-mode of the thermal polarized dust emission
@@
Figures 12 and 13 show the projected T-mode (temperature ) and B-mode (polarization) power spectra of the polarized dust emission for the same conditions. The spectral data were taken from Fig. 6 of the Ref. [39].
The dashed curves in Figs. 11,12, and 13 indicate the best fit by the stretched exponential spectral law Eq. (23), that means the importance of the mean magnetic field in this simulation. The same level of randomization of the magnetic field and the polarized emission confirms the imposition of the magnetic field level of randomization on the polarized emission (cf previous Section).
Figures 14 and 15 show the T-mode (temperature ) and B-mode (polarization) power spectra of the polarized dust emission computed using the GNILC-cleaned full-sky maps obtained using the Planck data at 353 GHz. The spectral data were taken from Fig. 11 of a paper [22]. The dashed curves in Figs. 14 and 15 indicate the best fit by the stretched exponential spectral law Eq. (23) (cf Figs. 12 and 13).
@@
In a recent paper [32] the third Planck public release’s maps were used to characterize the polarized dust emission at high Galactic latitudes. Using these maps the angular power spectra of the B-mode were computed at 353 GHz channel for six sky regions. The largest region covers 71 % of the sky (naturally, the best measurement results were obtained in this region). Figure 16 shows the power spectrum of the B-mode for this region. The spectral data were taken from the Table C.1 of the Ref. [32]. The dashed curve in Figs. 16 indicates the best fit by the stretched exponential spectral law Eq. (23).
@@ 7 Conclusions and Discussion
Analogously one can conclude that in the regions providing the main contribution to the all-sky map for the Faraday rotation measure the mean magnetic field also does not play an important role. Therefore the spectra corresponding to the magnetic field $B_{LoS}$ Fig. 7 and integrated electron density (electron dispersion measure) Fig. 8, obtained by disentangling the all-sky Faraday rotation map, are well fitted by Eq. (22).
The difference in the levels of randomization could be also partially related to the fact that for the smaller and middle frequencies the synchrotron and free-free emissions dominate the Galactic foreground whereas for the large frequencies the dust emission takes this role (see Fig. 5 of Ref. 14). The problem is that the observed angular power spectra of the synchrotron and free-free emission for the middle and high frequencies are rare in the literature as well as the spectra of the dust emission for small frequencies. Maybe the above consideration can be encouraging to solve this problem.
@@
Despite the vast differences in the values of physical parameters and spatio-temporal scales between the numerical simulations and the astrophysical observations, there is a quantitative agreement between the results of the astrophysical observations and the numerical simulations in the frames of the distributed chaos approach.

### v4

@@ 5 Observations in the Galactic magnetized plasma
In the paper (Hutschenreuter et al. 14) a disentangling of an all-sky map for average intragalactic $B_{LoS}$ from the Faraday effect map was reported. Some complementary tracers of the $n_{e}$ (the free-free map of the Planck survey 39, extra-Galactic Faraday data 55, a $H-\alpha$ map 9, pulsar data 28) were also used for the purpose. An all-sky map of the electron dispersion measure (the integrated electron density) was also constructed.
Figure 12 shows the power spectrum of $B_{LoS}$. The spectral data were taken from Figs. 9a of the paper (Hutschenreuter et al. 14). The dashed curve in Fig. 12 indicates the best fit by the stretched exponential spectral law Eq. (22).
An all-sky map of the integrated electron density (electron dispersion measure) was also obtained by this method. Figure 13 shows the power spectrum of the electron dispersion measure. The spectral data were taken from Figs. 9b of the paper (Hutschenreuter et al. 14). The dashed curve in Fig. 13 indicates the best fit by the same stretched exponential spectral law Eq. (22). Apparently the magnetic field imposes its degree of randomization (the $\beta=1/4$) on the electron dispersion measure.
@@ 6 Synchrotron emission
The astroparticles (relativistic electrons and positrons) produce the synchrotron emission moving through the magnetic field. Therefore, the diffuse (polarized) synchrotron emission is an effective tracer of the magnetic field in the magnetized non-thermal plasma. This emission is also one of the main components of the polarized foreground for the cosmic microwave background radiation. Therefore, studying its spectral properties is important for understanding the magnetized intragalactic plasma’s physical processes and obtaining clean CMB maps.
@@
In a paper [11] results obtained using this technique for the Parkes 2.4 GHz survey of synchrotron polarised emission for the Southern Galactic plane were reported. Figure 15 shows the power spectrum for the B-mode. The spectral data were taken from Fig. 3 of the Ref. (Giardino et al. 11). The dashed curve in Fig. 15 indicates the best fit by the stretched exponential spectral law Eq. (22).
@@ 7 Dust emission
Figure 17: As in Fig. 16. but for the T-mode of the thermal polarized dust emission
Figure 18: As in Fig 16 but for the B-mode of the thermal polarized dust emission
Polarized dust emission is not only one of the most reach sources of information about the Galactic magnetic field (see, for instance, a recent paper by 13, and references therein) but also the main obstacle for observing (detecting) the gravitational waves in primordial CMB B-modes, especially for the high-frequency Planck’s channels [21].
@@
Figures 17 and 18 show the projected T-mode (temperature ) and B-mode (polarization) power spectra of the polarized dust emission for the same conditions. The spectral data were taken from Fig. 6 of the Ref. [51].
The dashed curves in Figs. 16,17, and 18 indicate the best fit by the stretched exponential spectral law Eq. (23), that means the importance of the mean magnetic field in this simulation. The same level of randomization of the magnetic field and the polarized emission confirms the imposition of the magnetic field level of randomization on the polarized emission (cf previous Section).
Figures 19 and 20 show the T-mode (temperature ) and B-mode (polarization) power spectra of the polarized dust emission computed using the GNILC-cleaned full-sky maps obtained using the Planck data at 353 GHz. The spectral data were taken from Fig. 11 of a paper [29]. The dashed curves in Figs. 19 and 20 indicate the best fit by the stretched exponential spectral law Eq. (23) (cf Figs. 17 and 18).
@@
In a recent paper [40] the third Planck public release’s maps were used to characterize the polarized dust emission at high Galactic latitudes. Using these maps the angular power spectra of the B-mode were computed at 353 GHz channel for six sky regions. The largest region covers 71 % of the sky (naturally, the best measurement results were obtained in this region). Figure 21 shows the power spectrum of the B-mode for this region. The spectral data were taken from Table C.1 of the Ref. [40]. The dashed curve in Fig. 21 indicates the best fit by the stretched exponential spectral law Eq. (23).
@@ 8 Conclusions and Discussion
Analogously, one can conclude that in the regions providing the main contribution to the all-sky map for the Faraday rotation measure the mean magnetic field also does not play an important role. Therefore, the spectra corresponding to the magnetic field $B_{LoS}$ Fig. 12 and integrated electron density (electron dispersion measure) Fig. 13, obtained by disentangling the all-sky Faraday rotation map, are well fitted by Eq. (22).
The difference in the levels of randomization could also be partially related to the fact that for the smaller and middle frequencies the synchrotron and free-free emissions dominate the Galactic foreground whereas for the large frequencies the dust emission takes this role (see Fig. 5 of Ref. 21). The problem is that the observed angular power spectra of the synchrotron and free-free emission for the middle and high frequencies are rare in the literature as well as the spectra of the dust emission for small frequencies. Maybe the above consideration can be encouraging to solve this problem.
@@
The background radiation is considerably less randomized ($\beta=1/2$) than the Galactic foreground radiation ($\beta=1/4$, and $\beta=1/5$). This difference can be used for effective cleaning of the raw CMB data from the foreground.
Despite the vast differences in the values of physical parameters and spatio-temporal scales between the numerical simulations and the astrophysical observations, there is a quantitative agreement between the results of the astrophysical observations and the numerical simulations in the framework of the distributed chaos approach.